Hi,
Assuming that at least one coin is used, how many different sums of money can be made from the following coins: a penny, a nickel, a quarter, a fifty-cent piece, and a dollar?
I know the answer is 63 ways, but i have no idea how to approach the question and get the answer

Answer by sudhanshu_kmr(1152)


For each coin there is two ways, either include the coin or exclude it.
so, each coin can be selected in 2 ways, and total 5 coins
total no. of ways = 2*2*2*2*2 = 32
but there will one single case in which no coin selected, so discard it.
total no. of ways when at least one coin= 32-1 = 31

if any doubt, feel free to contact.